A sodium sulfate solution was electrolyzed using inert Pt electrodes. The cathode reaction was: 2 H2 0  +  2 e -  →  H2  + 2 OH-. If a current of 3.0 amp was used for 30 min, what weight of H2 gas would be produced? 


1
Expert's answer
2021-03-02T01:03:12-0500

Given, Current "i=3Amp."

Time "t=30min=30\times 60=1800s"


Cathode Reaction:

"2H_2O+2e^-\rightarrow H_2+2OH^-"


Then Amount of charge passes "Q=I\times t=3\times 1800" columbs.

The amount of chlorine liberated by passing "3\times 1800" coulomb of electric charge.  "=\dfrac{2}{2\times 96500}\times3\times 1800=0.05595 mole"



Volume of "H_2"  liberated at NTP "= 0.05595\times 22.4 = 1.2534 L"